OpenStudy (anonymous):

The equation, in general form, of the line that passes through the point (2,9) and is perpendicular to the line 4 x + 9 y + 1 = 0 is Ax + By + C = 0, where A = B = C =

hartnn (hartnn):

can you find slope of 4 x + 9 y + 1 = 0 ?

OpenStudy (anonymous):

ya i know how to fine the slop

hartnn (hartnn):

so whats the slope of that line ?

OpenStudy (anonymous):

isent it -4/9

hartnn (hartnn):

correct! now perpendicular line have product of slopes = -1

hartnn (hartnn):

so, m * (-4/9) = -1 m=... ?

OpenStudy (anonymous):

9/4?

hartnn (hartnn):

correct! m = 9/4 x1 =2 y1 =9 can you find the equation ?

hartnn (hartnn):

y-y1 = m (x-x1) slope point form

OpenStudy (anonymous):

it would be Y=9/4x+9/2

hartnn (hartnn):

yes, to eliminate the fraction, multiply 4 on both sides

OpenStudy (anonymous):

that would make it 4y=9x+18

OpenStudy (anonymous):

oh i got it thank you

hartnn (hartnn):

correct now to get that form, just subtract 4y from both isdes... ohh, welcome ^_^